# Primordial Black Hole stellar microlensing constraints: understanding their dependence on the density and velocity distributions

### v1

@@ 1 Introduction
In the 1990s the EROS [12], MACHO [13] and OGLE [14] collaborations began monitoring millions of stars in the Magellanic Clouds in order to probe COs in the Milky Way (MW) halo. The typical duration of a Large Magellanic Cloud (LMC) microlensing event caused by a CO with mass $M$ is $\sim(M/M)^{1/2}\,{\rm yr}$, see Eq. (2.2). Therefore a long-exposure survey is required to probe multi-solar mass COs, and a high-cadence survey to probe planetary mass COs. The EROS-2 survey [15] excluded COs with mass $10^{-7}\lesssim M/M_{\odot}\lesssim 10$ making up all of the DM in the MW halo. In recent years the upper end of the excluded mass range has been extended to $\sim 10^{3}M_{\odot}$ by combining data from EROS-2 and MACHO [16] and then to $\sim 10^{4}M_{\odot}$ by the OGLE-III and -IV long-exposure surveys [17, 18]. Very recently the minimum excluded mass has been reduced to $10^{-9}M_{\odot}$ by the OGLE high-cadence survey [19].
@@
Motivated by this apparent discrepancy, we revisit the dependence of the microlensing differential event, and constraints on the halo fraction in COs, on the DM distribution. In Sec. 2 we present the halos models that we use, their differential event rates and the exclusion limits on the halo fraction of COs from a toy long-exposure microlensing survey. In Sec. 3 we explore the origin of the behaviour of the microlensing differential event rate, and the accuracy of assuming a Maxwellian velocity distribution. Finally, we conclude with summary and discussion in Sec. 4
@@ 2.1 Differential event rate
$f({\bf v})\,{\rm d}^{3}{\bf v}=\frac{1}{\pi^{3/2}v_{\rm c}^{3}}\exp{\left(-\frac{v^{2}}{v_{\rm esc}^{2}}\right)}\,{\rm d}^{3}{\bf v}\,,$ (2.5)
@@
Evans’ power law (PL) models [33] are often used for studying the dependence of the microlensing differential rate on the DM density distribution (e.g. Refs. [34, 35, 13, 36, 20, 21]) as there are tractable expressions for the differential event rate, which include the (non-Maxwellian) velocity distribution (see Appendix B of Ref. [35]). These models are self-consistent, in that the velocity distribution is calculated from the potential of the halo. However close to the MW disc (i.e. at small line of sight distances $x$), the disc potential is non-negligible and will affect the velocity distribution, in particular increasing the velocity dispersion. The PL halo models have a density profile  [^2] We focus on the spherical case, $q=0$, in Eq. (2) of Ref. [33].
@@
The origin of the power law scalings of the rate of long- and short-duration events was explained by Mao and Paczysnski [38]. The duration of an event is $\hat{t}=R_{\rm E}/v_{\rm t}$, where $v_{\rm t}$ is the tangential velocity which is given by $v_{\rm t}=v\sin{i}$, where $v$ is the relative velocity between the lens and the observer and $i$ is the angle of inclination between the velocity and the line of sight. Long-duration events are caused by lenses which move almost along the light of sight. In this case $\hat{t}\propto\sin^{-1}{i}\propto{i}^{-1}$. The total number, $N$, (rather than differential rate) of such events is proportional to the solid angle, $(1-\cos{i})$, divided by $\hat{t}$, so that $N\propto i^{2}/i^{-1}\propto i^{3}\propto\hat{t}^{-3}$. Short-duration events are caused by lenses that are very close to either the observer ($x\sim 0$) or lens ($x\sim 1$) where the Einstein radius, $R_{\rm E}(x)$ is small. For lenses close to the observer $\hat{t}\propto R_{\rm E}\propto[xL]^{1/2}$ and the total number of these events is $N\propto R_{\rm E}xL/\hat{t}\propto\hat{t}^{3}$, with a similar argument leading to the same scaling for lenses close to the source.
@@ 2.2 Constraints
where $E$ is the exposure in star years and $\epsilon(\hat{t})$ is the detection efficiency i.e. the probability that a microlensing event that occurs with duration $\hat{t}$ is detected. For concreteness, we consider a toy long-exposure microlensing survey, which roughly mimics the OGLE 20-year survey [18]  [^3] Our conclusions would also apply, qualitatively, to a short-exposure, high-cadence microlensing survey that is sensitive to lighter CO, such as Ref. [19].. We assume that 79 million stars are observed for 7000 days, giving an exposure of $E=5.5\times 10^{11}$ star days. We use the efficiency for observing field number 501 from Ref. [39], which is shown, as a function of Einstein diameter crossing time, in Fig. 3. We calculate exclusion limits as for the ‘strict’ case in Ref. [17], assuming that none of the observed events are due to dark COs  [^4] If instead some of the observed events are allowed to be due to COs, the contribution of stars to the differential event rate has to be included in Eq. (2.9), and the exclusion limit on $f$ is weakened by a factor of a few for masses $10^{-4}\lesssim M/M_{\odot}\lesssim 10$.. For smoothly distributed COs the probability distribution of the number of observed events is Poissonian. Therefore, if no events are observed, the $95\%$ confidence limit on the fraction of the halo in COs is given by $f=3.0/N_{{\rm exp}}$.
@@ 3.1.1 Short-duration tail
$+\left.\int_{0}^{1/2}\tilde{\rho}(x)x^{2}(1-x)^{2}\exp{\left[-x(1-x)T\right]}\,{\rm d}x\right\}\,,$
@@
where in the final line we use the fact that the DM density at the LMC is much less than the local DM density, $\tilde{\rho}(1)\ll 1$. For the SH model Eq. (3.4) agrees with the full expression, Eq. (3.1), at the $5\%$ level (and at the $1\%$ level if the contribution from lens close to the source, $\tilde{\rho}(1)$, is included). The amplitude of the short-duration tail is determined by the product of the local dark matter density and the fourth power of the circular speed. A larger circular speed means the typical speed of the COs is larger. This increases the rate at which the COs enter the microlensing tube and also decreases the duration of the events they cause.
@@ 4 Summary
The microlensing differential event rate (i.e. the rate at which events of a given timescale occur), depends on the CO mass, density and velocity distributions. Uncertainties in the CO density and velocity distributions therefore lead to uncertainties in the constraints on the abundance of CO. We have focussed on understanding the dependence of the differential event rate, and hence the constraints, on the CO density and velocity distributions. We assume that the DM is smoothly distributed and use Evan’s PL halo models [33] as benchmarks, as they have tractable analytic expressions for the differential event rate which do not assume that the CO velocity distribution is Maxwellian [35].

### v2

@@ 1 Introduction
In the 1990s the EROS [12], MACHO [13] and OGLE [14] collaborations began monitoring millions of stars in the Magellanic Clouds in order to probe COs in the Milky Way (MW) halo. The typical duration of a Large Magellanic Cloud (LMC) microlensing event caused by a CO with mass $M$ is $\sim(M/M_{\odot})^{1/2}\,{\rm yr}$, see Eq. (2.2). Therefore a long-exposure survey is required to probe multi-solar mass COs, and a high-cadence survey to probe planetary mass COs. The EROS-2 survey [15] excluded COs with mass $10^{-7}\lesssim M/M_{\odot}\lesssim 10$ making up all of the DM in the MW halo. In recent years the upper end of the excluded mass range has been extended to $\sim 10^{3}M_{\odot}$ by combining data from EROS-2 and MACHO [16] and then to $\sim 10^{4}M_{\odot}$ by the OGLE-III and -IV long-exposure surveys [17, 18]. Very recently the minimum excluded mass has been reduced to $10^{-9}M_{\odot}$ by the OGLE high-cadence survey [19].
@@
Motivated by this apparent discrepancy, we revisit the dependence of the microlensing differential event, and constraints on the halo fraction in COs, on the DM distribution. Our aim is to understand how the constraints depend on the DM density and velocity distributions, in order to elucidate what is required for a reliable calculation of the constraints. In Sec. 2 we present the halo models that we use, their differential event rates and the exclusion limits on the halo fraction of COs from a toy long-exposure microlensing survey. In Sec. 3 we explore the origin of the behaviour of the microlensing differential event rate, and the accuracy of assuming a Maxwellian velocity distribution. Finally, we conclude with summary and discussion in Sec. 4
@@ 2.1 Differential event rate
$f({\bf v})\,{\rm d}^{3}{\bf v}=\frac{1}{\pi^{3/2}v_{\rm c}^{3}}\exp{\left(-\frac{v^{2}}{v_{\rm c}^{2}}\right)}\,{\rm d}^{3}{\bf v}\,,$ (2.5)
@@
Evans’ power law (PL) models [33] have tractable expressions for the differential event rate, which include the (non-Maxwellian) velocity distribution (see Appendix B of Ref. [35]). They are therefore often used as benchmarks for studying the dependence of the microlensing differential rate on the DM density distribution (e.g. Refs. [34, 35, 13, 36, 20, 21]). These models are self-consistent, in that the velocity distribution is calculated from the potential of the halo. However close to the MW disc (i.e. at small line of sight distances $x$), the disc potential is non-negligible and will affect the velocity distribution, in particular increasing the velocity dispersion. The PL halo models have a density profile  [^2] We focus on the spherical case, $q=0$, in Eq. (2) of Ref. [33].
@@
The origin of the power law scalings of the rate of long- and short-duration events was explained by Mao and Paczysnski [38]. The duration of an event is $\hat{t}=2R_{\rm E}/v_{\rm t}$, where $v_{\rm t}$ is the tangential velocity which is given by $v_{\rm t}=v\sin{i}$, where $v$ is the relative velocity between the lens and the observer and $i$ is the angle of inclination between the velocity and the line of sight. Long-duration events are caused by lenses which move almost along the light of sight. In this case $\hat{t}\propto\sin^{-1}{i}\propto{i}^{-1}$. The total number, $N$, (rather than differential rate) of such events is proportional to the solid angle, $(1-\cos{i})$, divided by $\hat{t}$, so that $N\propto i^{2}/i^{-1}\propto i^{3}\propto\hat{t}^{-3}$. Short-duration events are caused by lenses that are very close to either the observer ($x\sim 0$) or lens ($x\sim 1$) where the Einstein radius, $R_{\rm E}(x)$ is small. For lenses close to the observer $\hat{t}\propto R_{\rm E}\propto[xL]^{1/2}$ and the total number of these events is $N\propto R_{\rm E}xL/\hat{t}\propto\hat{t}^{3}$, with a similar argument leading to the same scaling for lenses close to the source.
@@ 2.2 Constraints
where $E$ is the exposure in star years and $\epsilon(\hat{t})$ is the detection efficiency i.e. the probability that a microlensing event that occurs with duration $\hat{t}$ is detected. For concreteness, we consider a toy long-exposure microlensing survey, which roughly mimics the OGLE 20-year survey [18]  [^3] Our conclusions would also apply, qualitatively, to a short-exposure, high-cadence microlensing survey that is sensitive to lighter CO, such as Ref. [19].. We assume that 79 million stars are observed for 7000 days, giving an exposure of $E=5.5\times 10^{11}$ star days. We use the efficiency for observing field number 501 from Ref. [39], which is shown, as a function of Einstein diameter crossing time, in Fig. 3. We calculate exclusion limits as for the ‘strict’ case in Ref. [17], assuming that none of the observed events are due to dark COs  [^4] If instead some of the observed events are allowed to be due to COs, the contribution of stars to the differential event rate has to be included, and the exclusion limit on $f$ is weakened by a factor of a few for masses $10^{-4}\lesssim M/M_{\odot}\lesssim 10$.. For smoothly distributed COs the probability distribution of the number of observed events is Poissonian. Therefore, if no events are observed, the $95\%$ confidence limit on the fraction of the halo in COs is given by $f=3.0/N_{{\rm exp}}$.
@@ 3.1.1 Short-duration tail
$+\left.\int_{1/2}^{1}\tilde{\rho}(x)x^{2}(1-x)^{2}\exp{\left[-x(1-x)T\right]}\,{\rm d}x\right\}\,,$
@@
where in the final line we use the fact that the DM density at the LMC is much less than the local DM density, $\tilde{\rho}(1)\ll 1$. For the SH model Eq. (3.4) agrees with the full expression, Eq. (3.1), at the $5\%$ level (and at the $1\%$ level if the contribution from lenses close to the source, $\tilde{\rho}(1)$, is included). The amplitude of the short-duration tail is determined by the product of the local dark matter density and the fourth power of the circular speed. A larger circular speed means the typical speed of the COs is larger. This increases the rate at which the COs enter the microlensing tube and also decreases the duration of the events they cause.
@@ 4 Summary
The microlensing differential event rate (i.e. the rate at which events of a given timescale occur), depends on the CO mass, density and velocity distributions. Uncertainties in the CO density and velocity distributions therefore lead to uncertainties in the constraints on the abundance of CO. We have focussed on understanding the dependence of the differential event rate, and hence the constraints, on the CO density and velocity distributions. This is a necessary step towards an accurate calculation of the constraints. We assume that the DM is smoothly distributed and use Evan’s PL halo models [33] as benchmarks, as they have tractable analytic expressions for the differential event rate which do not assume that the CO velocity distribution is Maxwellian [35].